OpenStudy (anonymous):

How much would $400 invested at 9% interest compounded continuously be worth after 3 years? Round your answer to the nearest cent. A. $475.73 B. $508.00 C. $406.60 D. $523.99

OpenStudy (danjs):

Do you remember the continuously compounded interest formula from the previous problems?

OpenStudy (anonymous):

this one is A(t)=p *e^rt

OpenStudy (danjs):

right on, p is the principle amount invested 400 r is the rate, 0.09 t is the time 3 yrs

OpenStudy (here_to_help15):

This is really simple, you are just plugging known values into the formula A(t) = Pe^(rt) P = $400 r = 0.09 t = 3 (years) A(3) = 400e^(0.09*3) A(3) = $523.99 *I rounded to the nearest penny, which is customary when rounding dollar amounts.

OpenStudy (here_to_help15):

@DanJS why have you blocked me?